OpenStudy (anonymous):

A company distributes free candies to all the students of x schools. Each school has (x + 1) classes. The number of students in each class is 3 more than the number of classes in each school. Each student is given 4 candies. How can you calculate the total number of students in each school?

OpenStudy (anonymous):

So we have X schools. Each school has X+1 classes. So we have \[x(x+1)=x^{2}+x\] classes. Each class has 3 more students than the number of classes so: \[x^{2}+x+3\] students in each class for a total of \[(x^{2}+x)classes \times (x^{2}+x+3) \frac{students}{class}\] students. This gives us:\[x^{4}+x^{3}+3x^{2}+x^{3}+x^{2}+3x=x^{4}+2x^{3}+4x^{2}+3x\] students.

OpenStudy (anonymous):

@twvogels can you help me with anther question

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Greg drew 3 squares with each side equal to z units. For each square, he does something different to it according to each part below. Greg decreased the length and width of the square by t units each. What will be the change in the area of the square? Show your work

OpenStudy (anonymous):

@twvogels

OpenStudy (anonymous):

 So for the large square the area is: \[A_L=z^{2}\] and for the small one it's: \[A_S=(z-t)^{2}\] Then we can subtract to get the difference. \[A_L−A_S\] \[z^{2}-(z-t)^{2}\]\[z^{2}-(z^{2}-2zt+t^{2})\]\[z^{2}-z^{2}+2zt-t^{2})\]\[2zt-t^{2}\]

OpenStudy (anonymous):

WOW thank you so much you are amazing one more question if that is ok with you

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

A: Divide (10x^4y^3 + 5x^3y^2 - 15x^2y - 25x^2y^4) by -5x^2y. Show your work. B: How would your answer in Part A be affected if the x^2 variable in the denominator was just an x? C:What is the degree and classification of the polynomial you got in Part A? @twvogels

OpenStudy (anonymous):

\[\frac{10x^{4}y^{3}+5x^{3}y^{2}-15x^{2}y-25x^{2}y^{4}}{-5x^{2}y}\] Just to make sure I read it right.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, this is going to take some time. Hold on.

OpenStudy (anonymous):

\[-(2x^{2}y^{2}+xy-5y^{3}-3)\]

OpenStudy (anonymous):

wait which part is that the answer for is it for A

OpenStudy (anonymous):

Yes, that's Part A.

OpenStudy (anonymous):

Can you give me the answers for the other parts too

OpenStudy (anonymous):

Ok, I'll need to work them out. Give me a minute.

OpenStudy (anonymous):

B is just the same as a, multiplied by the -5 we removed: \[5(2x^{2}y^{2}+xy−5y^{3}−3)\]

OpenStudy (anonymous):

so \[5(2x^2y^2+xy-5y^3-3)\] is the answer

OpenStudy (anonymous):

to part b

OpenStudy (anonymous):

so the answer for c would be 3rd degree and a classification of 4 right @twvogels

OpenStudy (anonymous):

Yes that's correct.

OpenStudy (anonymous):

ok thank you so much i became a fan